OpenStudy (anonymous):

Select the number property that will justify the step in the following expression. 6 + (14 + 70) = (6 + 14) + 70 = Associative-Addition, Transitive, Commutative-Addition, Identity-Addition 20 + 70 = 90

OpenStudy (anonymous):

it is the associative law of addition

OpenStudy (anonymous):

Continue To Help?

OpenStudy (anonymous):

the associative law says \[a+(b+c)=(a+b)+c\]

OpenStudy (anonymous):

it is easy to remember which one is the associative law: the parentheses move, but the letters stay in the same order do you have another question you need help with?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

shoot, if i can help i will

OpenStudy (anonymous):

Select the number property that will justify the step in the following expression. 2(5 17) = (2 5) 17 = Associative-Addition, Commutative-Multiplication, Identity-Addition, Associative-Multiplication 10 17 = 170

OpenStudy (anonymous):

i take it this is \[2\times (5\times 17)=(2\times 5)\times 17\]

OpenStudy (anonymous):

as in the previous one, the numbers stay in the same order, but the parentheses move this is the associative law, but this time for multiplication rather than addition

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

yw